# A Hypernetwork-Based Approach to KAN Representation of Audio Signals

### v1

@@
Abstract
Implicit neural representations (INR) have gained prominence for efficiently encoding multimedia data, yet their applications in audio signals remain limited. This study introduces the Kolmogorov-Arnold Network (KAN), a novel architecture using learnable activation functions, as an effective INR model for audio representation. KAN demonstrates superior perceptual performance over previous INRs, achieving the lowest Log-Spectral Distance of 1.29 and the highest Perceptual Evaluation of Speech Quality of 3.57 for 1.5 s audio. To extend KAN’s utility, we propose FewSound [^1] The source code can be accessed at https://github.com/gmum/fewsound.git, a hypernetwork-based architecture that enhances INR parameter updates. FewSound outperforms the state-of-the-art HyperSound, with a 33.3% improvement in MSE and 60.87% in SI-SNR. These results show KAN as a robust and adaptable audio representation with the potential for scalability and integration into various hypernetwork frameworks.
@@ Abstract
1 Introduction
Implicit neural representations (INRs) represent discrete data in neural network weights. INR finds many applications in the case of image (Klocek et al., 2019), video (Chen et al., 2022), or 3D objects (Spurek et al., 2020). In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data. HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes on raw input sound and produces weight for INR-based sound representation. INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes. Siamese SIREN (Lanzendörfer & Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
@@ 1 Introduction
This paper demonstrates that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm. For single audio INR, KAN trains more rapidly and achieves better results. Setting up the conditions for a KAN is not immediately straightforward. We show that hypernetworks can generate KANs. In practice, we can directly produce KAN parameters or use updates for the universal KAN-based architecture. Therefore, we propose two methods. The first is a classical HyperSound (Szatkowski et al., 2023) with a KAN target network. The second uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR, which is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017). Since in a few-shot approach, there does not exist a model dedicated to audio signals, we introduce FewSound. FewSound takes raw sound on input and uses hypernetwork to generate updates to universal weights of INR dedicated to encoding sound. In our work, we examine six approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a). In the model, we simultaneously train the hypernetwork weights and the universal weights. Thus, two settings show that KAN can be effectively modeled by hypernetworks.
Our study demonstrates that KAN model can be viewed as INR of sound. Such models yield favorable outcomes when dealing with individual audio signals and excel in large Hypernetwork-based models trained on large datasets.
@@
  • showing KAN’s efficacy for continuous sound representation,
@@
  • analysis of the parameters’s influence on the efficacy of the proposed method.
2 Related Work
@@
Implicit neural representations (INR) are data approximations based on neural networks relying on coordinate inputs. They enable discrete information such as pixels (Klocek et al., 2019; Chen et al., 2022), cloud points (Spurek et al., 2020), or audio magnitude information (Szatkowski et al., 2023) to be represented as continuous functions. The separation of the representation from spatial resolution makes the representation more memory efficient. Such information is limited by the networks’ architecture instead of the grid resolution. INRs are used in many signal processing problems, including 3D reconstruction (Mildenhall et al., 2020), super-resolution (Chen et al., 2022), compression (Strümpler et al., 2022; Lanzendörfer & Wattenhofer, 2023) or audio scene representation (Su et al., 2022).
@@
NERF (Neural Radiance Field) (Mildenhall et al., 2020) is a type of network originally designed for scene representation. Based on the 5D input of spatial location $(x,\,y,\,z)$ and viewing direction ($\theta,\,\phi$), the MLP network outputs color c = $(r,\,g,\,b)$ and volume density ($\sigma$). This allows the generation of 3D reconstructions of scenes from a set of sparse views.
@@
SNAC (Siuzdak, 2024) uses the encoder based on the hierarchical tokens like the previous approach. The distinction is that the coarse tokens are sampled less frequently, resulting in covering broader time windows. SNAC not only decreases the bitrate required for reconstruction, but also provides useful information in language modeling tasks.
@@
We consider six variants of INRs: NeRF, SIREN, RFF, FINER, WIRE and KAN. We use NeRF to refer to a network with positional encoding, followed by a MLP with ReLU activation. The positional encoding, described by $\gamma(t)=\left[\sin(2^{0}\pi t),\cos(2^{0}\pi t),\ldots,\sin(2^{L-1}\pi t),\cos(2^{L-1}\pi t)\right]$ is taken from the original NeRF (Mildenhall et al., 2020).
SIREN (Sitzmann et al., 2020) is a MLP with sine activation: $a(x)=\sin(\omega_{0}x)$, where $\omega_{0}$ is a hyperparameter.
RFF (Random Fourier Features) (Tancik et al., 2020) is used to refer to a multilayer perceptron with ReLU activation and input embedding defined as $\gamma(t)=\left[\cos(2\pi\mathbf{B}t),\sin(2\pi\mathbf{B}t)\right],$ where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is a random projection matrix, sampled from $\mathcal{N}(0,\sigma^{2})$. In our case $d$, the dimensionality of the input, is always equal to 1, while $\sigma$ and $m$ are hyperparameters.
WIRE (Saragadam et al., 2023) is a MLP with complex Gabor wavelet activation $a(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$, where $\omega_{0}$ and $s_{0}$ are hyperparameters.
FINER (Liu et al., 2024b) is a MLP with activation function defined as $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$, where $\omega_{0}$ is a hyperparameter. The bias of its first layer is initialized by sampling from $\mathcal{U}(-k,k)$, where $k$ is a hyperparameter.
Our Kolmogorov-Arnold network implementation consists of the same positional encoding as in NeRF, followed by KAN layers. Each KAN layer can be described by $\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}spline(x),$ where $w_{b}$ and $w_{s}$ are matrices of learnable weights, and $spline(x)$ is a linear combination of B-splines such that $spline(x)=\sum_{i}{c_{i}B_{i}(x)},$ with learnable $c_{i}$s. It is parametrized by spline order – the degree of the used spline, and grid size – the number of knots defining the spline.
3 KAN in hypernetwork-based architectures
@@
Figure 1: The overview of FewSound architecture. The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights. The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$. The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
Table 1: Reconstruction metrics for 1.5 s and 3 s audio. We report mean and standard deviation across 110 samples, each of a different speaker of VCTK dataset (Veaux et al., 2013).
@@
Few-shot learning for INR
The terminology of few-shot learning is often inconsistent due to varying definitions in existing research. For a standardized framework (Chen et al., 2019; Wang et al., 2020). However, such a nomenclature is dedicated to a classification, regression, or generative model task. In the case of the INR-based model, we refer to (Tancik et al., 2021). We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
FewSound
The parameters for the final target model are calculated with the formula in Eq. 1.
$\theta^{\prime}=\theta+\Delta_{\theta}=\theta+H(\mathbf{E}_{\mathcal{S}},\,\mathbf{E}_{\theta}).$ (1)
Training
For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$. The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$. During training, we first sample the batch of sounds $\mathcal{B}$ from the given dataset $X$. Next, for each task data, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1). Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
$\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$ (2)
4 Experiments
In this section, we perform an experimental analysis. We divide it into two subsections: analysis of the performance of different INR architectures and experimental results of our hypernetwork approach, FewSound.
@@ 4 Experiments
4.1 Implicit Neural Representations
We compare the aforementioned INR architectures using two subsets of the VCTK (Veaux et al., 2013) dataset. Each of the subsets is generated by randomly choosing one recording per speaker, resampling it to 22.05 kHz and randomly cropping it to a desired length. We crop the recordings to 32,768 samples (about 1.5 seconds) in the first subset and to 65,536 (about 3 seconds) in the second subset. Both subsets consist of a total of 110 recordings. To make a fair comparison, all of the MLP based INRs have the same number of layers and the same sizes of layers, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models. All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov & Hutter, 2019) and the same loss function as in FewSound. Table 1 shows the results of the experiments. All INRs were selected to be close in the number of parameters. One may note that, with the exception of MSE, the shorter audios result in better quality. KAN is competitive for most metrics, achieving the best mean score in LSD, PESQ, STOI and CDPAM and second best PSNR for 1.5 s audio; as well as the best mean LSD, PESQ, STOI, and second best MSE, PSNR, SI-SNR, CDPAM for 3 s audio.
4.2 Hypernetworks
@@
When using hypernetworks, we only consider NeRF, SIREN, and KAN as the target network architectures. We compare our method with the state-of-the-art model HyperSound (Szatkowski et al., 2023) and follow the same experimental setting. We use the VCTK dataset resampled to $22,050$ Hz. The inputs have a length of 32,768 samples and are obtained by cropping the original recordings. The last 10 speakers are used as the validation set. All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch. We use AdamW optimizer (Loshchilov & Hutter, 2019) and OneCycleLR scheduler (Smith & Topin, 2018). The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is set to $1\mathrm{e}{-6}$.
@@ Results
Table 2: FewSound performance compared to HyperSound.
@@
Table 3: Performance of FewSound with different target network architectures.
MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
Table 4: Performance of HyperSound with different target network architectures. Results for SIREN and NERF are taken from the original paper, while we train the model using the KAN target network architecture ourselves. A target network with a significantly lower number of parameters (19K) is used in this case, as it achieves better results.
MSE LSD SI-SNR PESQ STOI CDPAM
@@
Ablation Study
Table 5 shows the impact of using different encoder architectures, mentioned in Section 2. SoundStream, SNAC, and Spectrogram Encoder were trained from scratch. In the case of Whisper, we base on the pre-trained checkpoints - tiny.en and large-v2 and finetune using the LoRA approach (Hu et al., 2022).
Table 5: FewSound performance with different encoder architectures.
@@
Table 6: KAN performance on the 1.5 s samples, with respect to the encoding length. The results are averaged across 10 samples.

### v2

@@
Abstract.
Implicit neural representations (INR) have gained prominence for efficiently encoding multimedia data, yet their applications in audio signals remain limited. This study introduces the Kolmogorov-Arnold Network (KAN), a novel architecture using learnable activation functions, as an effective INR model for audio representation. KAN demonstrates superior perceptual performance over previous INRs, achieving the lowest Log-Spectral Distance of 1.29 and the highest Perceptual Evaluation of Speech Quality of 3.57 for 1.5 s audio. To extend KAN’s utility, we propose FewSound, a hypernetwork-based architecture that enhances INR parameter updates. FewSound outperforms the state-of-the-art HyperSound, with a 33.3% improvement in MSE and 60.87% in SI-SNR. These results show KAN as a robust and adaptable audio representation with the potential for scalability and integration into various hypernetwork frameworks.
@@ Abstract.
conference: ; isbn: ;ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Learning paradigmsccs: Applied computing Sound and music computing
Figure 1. Sound representation is a challenging task. In classical analysis, we usually apply the FFT (Fast Fourier Transform), which analyzes frequencies by separating them into sinusoidal components. We propose to analyze sounds using KAN (Kolmogorov-Arnold Network), which uses trainable activation functions instead of classical layers. Such an operation can be seen as approximating signals by a linear combination of basis functions, bearing a similarity to the FFT or wavelets.
1. Introduction
Implicit neural representations (INRs) depict discrete data in neural network weights. INR finds many applications in the case of image (Klocek et al., 2019), video (Chen et al., 2022), or 3D objects (Spurek et al., 2020). In the case of sound, INRs are unexplored, and only a few approaches are dedicated to such data. HyperSound (Szatkowski et al., 2023) presents an autoencoder-based architecture that takes raw input sound and produces weight for INR-based sound representation. INRAS (Su et al., 2022) is an INR network that, by mapping the scene coordinates to the corresponding impulse responses, enables the modeling of spatial audio of interior scenes. Siamese SIREN (Lanzendörfer and Wattenhofer, 2023), is the extension of SIREN (Sitzmann et al., 2020) used for audio compression.
@@ 1. Introduction
This paper shows that KAN is effective in modeling individual audio signals and can be extended to various functions through the hypernetwork paradigm. For single audio INR, KAN trains more rapidly and achieves better results. Setting up the conditions for a KAN is not immediately straightforward. We show that hypernetworks can generate KANs. In practice, we can directly produce KAN parameters or use updates for the universal KAN-based architecture. Hence, we propose two methods, a classical HyperSound (Szatkowski et al., 2023) with a KAN target network. The second uses a few-shot setting (Tancik et al., 2021; Batorski et al., 2024) in INR, which is an implementation of the Model-Agnostic Meta-Learning algorithm (MAML) (Finn et al., 2017). As there is no few-shot model for audio signals, we introduce FewSound  [^1] The source code for the proposed method can be found at the following link: https://github.com/gmum/fewsound. It takes raw sound on input and uses a hypernetwork to generate updates to the universal weights of INR dedicated to encoding sound. In our work, we examine six approaches: SIREN (Sitzmann et al., 2020), NeRF (Mildenhall et al., 2020), Random Fourier Features (Tancik et al., 2020), WIRE (Saragadam et al., 2023), FINER (Liu et al., 2024b) and KAN (Liu et al., 2024a). In the model, we simultaneously train the hypernetwork weights and the universal weights. Thus, two settings show that KAN can be effectively modeled by hypernetworks.
Our study demonstrates that KAN can be viewed as INR of sound. Such models yield favorable outcomes when dealing with individual audio signals and excel in large Hypernetwork-based models trained on large datasets. We use KANs due to their spline-based activations. Since any function f (e.g., an audio representation of the magnitude over time) can be approximated by splines with $\sup\|f\_spline\|=\mathrm{O}(h^{4})$ (Hall and Meyer, 1976), KANs inherit this theoretical precision as a linear combination of splines, resulting in a polynomial bound. This suggests that KANs could be effective for audio reconstruction.
@@
  • introducing KAN-based sound representation and showing its efficacy,
@@
  • analysis of the parameters’ influence on the efficacy of the proposed method.
2. Related Work
@@
Implicit neural representations (INR) are data approximations based on neural networks relying on coordinate inputs. They enable discrete information such as pixels (Klocek et al., 2019; Chen et al., 2022), cloud points (Spurek et al., 2020), or audio magnitude information (Szatkowski et al., 2023) to be represented as continuous functions. The separation of the representation from spatial resolution makes the representation more memory efficient. Such information is limited by the networks’ architecture instead of the grid resolution. INRs are used in many signal processing problems, including 3D reconstruction (Mildenhall et al., 2020), super-resolution (Chen et al., 2022), compression (Strümpler et al., 2022; Lanzendörfer and Wattenhofer, 2023) or audio scene representation (Su et al., 2022).
@@
NeRF (Neural Radiance Field) (Mildenhall et al., 2020) models scenes via a 5D input—spatial location $(x,\,y,\,z)$ and viewing direction $(\theta,\,\phi)$—with an MLP outputting color $c=(r,\,g,\,b)$ and density $\sigma$, enabling 3D reconstruction from sparse views.
Figure 2. The overview of FewSound architecture. The input row sound is processed by encoding network $E(\cdot)$ and delivered to the fully connected Hypernetwork $H(\cdot)$ together universal weights. The Hypernetwork transforms them and returns the update of weights $\Delta\theta$ for the target INR $f_{\theta^{\prime}}$. The target network $f_{\theta^{\prime}}$ transforms time samples into the amplitude of the sound, dedicated to the considered task.
@@
SNAC (Siuzdak et al., 2024) employs a hierarchical token-based encoder, similar to prior work, but samples coarse tokens less frequently to span longer time windows. This reduces reconstruction bitrate and aids language modeling tasks.
@@
We consider six variants of INRs: NeRF, SIREN, RFF, FINER, WIRE and KAN. We use NeRF to refer to a network with positional encoding, followed by an MLP with ReLU activation. The positional encoding, described by
$\gamma(t)=\left[\sin(2^{0}\pi t),\cos(2^{0}\pi t),\ldots,\sin(2^{L-1}\pi t),\cos(2^{L-1}\pi t)\right]$
is taken from the original NeRF (Mildenhall et al., 2020).
SIREN (Sitzmann et al., 2020) is an MLP with sine activation: $a(x)=\sin(\omega_{0}x)$, where $\omega_{0}$ is a hyperparameter.
RFF (Random Fourier Features) (Tancik et al., 2020) is used to refer to an MLP with ReLU activation and input embedding defined as
$\gamma(t)=\left[\cos(2\pi\mathbf{B}t),\sin(2\pi\mathbf{B}t)\right],$
where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is a random projection matrix, sampled from $\mathcal{N}(0,\sigma^{2})$. In our case $d$, the dimensionality of the input, is always equal to 1, while $\sigma$ and $m$ are hyperparameters.
WIRE (Saragadam et al., 2023) is an MLP with complex Gabor wavelet activation $a(x)=e^{i\omega_{0}x}e^{-\lvert s_{0}x\rvert^{2}}$, where $\omega_{0}$ and $s_{0}$ are hyperparameters.
FINER (Liu et al., 2024b) is an MLP with activation function defined as $a(x)=\sin(\omega_{0}\lvert x+1\rvert x)$, where $\omega_{0}$ is a hyperparameter. The bias of its first layer is initialized by sampling from $\mathcal{U}(-k,k)$, where $k$ is a hyperparameter.
Our Kolmogorov-Arnold network implementation consists of the same positional encoding as in NeRF, followed by KAN layers. Each KAN layer can be described by $\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}spline(x),$ where $w_{b}$ and $w_{s}$ are matrices of learnable weights, and $spline(x)$ is a linear combination of B-splines such that $spline(x)=\sum_{i}{c_{i}B_{i}(x)}$ with learnable $c_{i}$s. It is parametrized by spline order – the degree of the spline, and grid size – the number of knots defining the spline.
3. KAN in hypernetwork-based architectures
@@
Few-shot learning for INR
The terminology of few-shot learning is often inconsistent due to varying definitions in existing research. For a standardized framework, see  (Chen et al., 2019; Wang et al., 2020). However, such a nomenclature is dedicated to a classification, regression, or generative model task. In the case of the INR-based model, we refer to (Tancik et al., 2021). We have a data set $X$ that contains sound samples, and our goal is to produce INR $f_{\theta}$ that encodes sound in network weights.
FewSound
Target model parameters are computed via Eq. 1.
(1) $\theta^{\prime}=\theta+\Delta_{\theta}=\theta+H(\mathbf{E}_{\mathcal{S}},\,\mathbf{E}_{\theta}).$
While FewSound accepts a fixed-size input (in our case 32768), it can be easily used on audio samples of arbitrary length by reconstructing them piece by piece and smoothing during merging.
Training
For training the model, we assume that the encoder $E(\cdot)$ is parametrized by $\gamma$, $E:=E_{\gamma}$. The universal weight encoder $G(\cdot)$ is parameterized by $\delta$, and the Hypernetwork $H(\cdot)$ by $\eta$, $H:=H_{\eta}$. During training, we first sample the batch of sounds $\mathcal{B}$ from the dataset $X$. Next, for each task, we calculate the update $\Delta_{\theta_{i}}$ using the hypernetwork and provide the updated parameters $\theta^{\prime}$ according to the rule given in (1). Finally, the objective to train the parameters of the system is calculated using the timesteps from [1, 1]11[-1,\,1][ - 1 , 1 ]:
(2) $\mathcal{L}_{HyperMAML}(f_{\theta})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta^{\prime}})=\sum_{\mathcal{T}_{i}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}_{i}}(f_{\theta+\Delta_{\theta}})~{},$
Table 1. Reconstruction metrics for 1.5 s and 3 s audio. We report mean and standard deviation across 110 samples, each of a different speaker of VCTK dataset (Veaux et al., 2013).
@@
4. Experiments
This section presents an experimental analysis in two parts: performance comparison of INR architectures and results of our hypernetwork method, FewSound.
@@ 4. Experiments
4.1. Implicit Neural Representations
We compare the aforementioned INR architectures using two subsets of the VCTK (Veaux et al., 2013) dataset. Each of the subsets is generated by randomly choosing one recording per speaker, resampling it to 22.05 kHz and randomly cropping it to a desired length. We crop the recordings to 32,768 samples (about 1.5 seconds) in the first subset and to 65,536 (about 3 seconds) in the second subset. Both subsets consist of a total of 110 recordings. To make a fair comparison, all of the MLP based INRs have the same number of layers and with the same sizes, while KANs have their architecture chosen in a way that the number of parameters is similar to the other models. All architectures are trained for 10,000 steps, using the AdamW optimizer (Loshchilov and Hutter, 2019) and the same loss function as in FewSound. Table 1 shows the results of the experiments. One may note that, with the exception of MSE, the shorter audios result in better quality. KAN is competitive for most metrics, achieving the best mean score in LSD, PESQ, STOI and CDPAM and second best PSNR for 1.5 s audio; as well as the best mean LSD, PESQ, STOI, and second best MSE, PSNR, SI-SNR, CDPAM for 3 s audio.
Further comparison of these INRs on additional datasets is available in section A.2.
4.2. Hypernetworks
@@
When using hypernetworks, we only consider NeRF, SIREN, and KAN as the target network architectures. We compare our method with the state-of-the-art model HyperSound (Szatkowski et al., 2023) and follow the same experimental setting. We use the VCTK dataset resampled to $22,050$ Hz. The inputs are truncated to length of 32,768 samples. The last 10 speakers are used as the validation set. All models are trained for 2,500 epochs, with 10,000 training samples chosen randomly for each epoch. We use AdamW optimizer (Loshchilov and Hutter, 2019) and OneCycleLR scheduler (Smith and Topin, 2018). The initial learning rate is set to $1\mathrm{e}{-5}$, with the exception of runs using the SIREN target network, where it is $1\mathrm{e}{-6}$.
@@ Results
Table 2. FewSound performance compared to HyperSound.
@@
Table 3. Effectiveness of FewSound with different target network architectures.
Target MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
Table 4. Performance of HyperSound with different target network architectures. Results for SIREN and NERF are taken from the original paper, while we train the model using the KAN target network architecture ourselves. A target network with a significantly lower number of parameters (19K) is used in this case, as it achieves better results.
Target MSE LSD SI-SNR PESQ STOI CDPAM
Network $\rightarrow 0$ $\rightarrow 0$ $\rightarrow 100$ $\rightarrow 4.5$ $\rightarrow 1$ $\rightarrow 0$
@@
Influence of encoder architecture
Table 5 shows the impact of using different encoder architectures, mentioned in Section 2. SoundStream, SNAC, and Spectrogram Encoder are trained from scratch. In the case of Whisper, we base on the pre-trained checkpoints - tiny.en and large-v2 and finetune using the LoRA approach (Hu et al., 2022). Among the encoders considered, SoundStream achieves the best results in terms of distance-based metrics such as PSNR, while Spectrogram Encoder excels in perceptual ones. SNAC doesn’t yield poor results, but it is essentially a less effective version of SoundStream. The Whisper models, by contrast, show significantly lower PSNR along with weaker PESQ score.
Table 5. FewSound results with different encoders.
@@
Effect of dataset variability
We extend the results for our FewSound model by training it on three additional speech datasets - noisy VCTK (Valentini-Botinhao et al., 2016), a version of the VCTK dataset with added noise, LibriSpeech (Panayotov et al., 2015) and LJ Speech (Ito and Johnson, 2017). Moreover, it is also evaluated on GTZAN (Tzanetakis and Cook, 2002) and UrbanSound8K (Salamon et al., 2014), which represent music and environmental noise, respectively. We follow the same setup as in our main experiment, with one exception—in this experiment, we use a hypernetwork MLP with 41.2M parameters instead of the standard one. In case of Librispeech, we train on the train-clean-100 partition and use the test-other partition as the validation set. Since PESQ, STOI and CDPAM are suitable only for clean speech datasets, they are omitted from the FewSound model assessment on noisy VCTK, GTZAN and UrbanSound8K. All obtained results are presented in Table 6. The results indicate that FewSound achieves lower MSE and higher SI-SNR on speech datasets than on music or environmental noise, likely due to encoder limitations in handling non-speech signals. Between the KAN and NeRF targets, KAN consistently outperforms in LSD and PESQ across all datasets. For SI-SNR and PSNR, performance is comparable, e.g., NeRF excels on LibriSpeech, while KAN leads on LJ Speech.
Table 6. FewSound performance across different datasets.
Dataset Target MSE PSNR LSD SI-SNR PESQ STOI CDPAM
[…]
VCTK NERF 0.006 28.39 1.50 5.03 1.41 0.79 0.30
KAN 0.006 28.35 1.46 5.11 1.45 0.79 0.34
noisy NERF 0.013 24.76 1.42 2.15 - - -
VCTK KAN 0.013 24.69 1.31 3.31 - - -
Libri NERF 0.016 23.8 2.42 -2.18 1.26 0.71 0.34
Speech KAN 0.018 23.47 1.93 -3.21 1.31 0.73 0.31
LJ NERF 0.013 24.80 1.66 -0.72 1.22 0.74 0.38
Speech KAN 0.012 25.11 1.51 -0.43 1.24 0.76 0.35
GTZAN NERF 0.031 21.21 1.49 -4.23 - - -
KAN 0.028 21.59 1.34 -3.50 - - -
Urban NERF 0.051 18.94 1.28 -12.93 - - -
Sound-8k KAN 0.049 19.15 1.22 -11.86 - - -
4.3. KAN Ablation Study
Table 7. KAN performance on the 1.5 s samples, with respect to the encoding length. The results are averaged across 10 samples.